f(h(x))= when f(x) = 2x-1 and h = x+4
I took (2x-1)(x+4) and that gave me 2x^+8x-x-4 which i 2x^+7x-4 but my options for an answer are 2x+3, 2x+7, 7, and 2x^+7x-8 - i dont get any of them and i have been working on this for 2 hours. Can anyone please help me - thank you.

Answer by scott8148(6628)

the notation is the thing

f(h(x)) means put h(x) in place of x in the function f(x)

f(h(x)) = 2(h(x)) - 1 = 2(x + 4) - 1 = 2x + 7